the ammonia evolved from the treatment of 0.3g of an organic compound for the estimation of nitrogen was passed in 100ml of 0.1 M sulphuric acid. The excess of acid required 20ml of 0.5 M sodium h hydroxide solution for complete neutralization. The organic compound is a) acetamide b) thiourea c) urea d) benzamide

Grade:12th pass

1 Answers

Vikas TU
7 years ago
NaOH mmoles that reacted with excess acid = 20*0.5 = 10 mm
excess acid mm => {from mole 2 conceptt}
moles x valence factor = cinstant.
Therefore,
nH2SO4 x 2 = 10 x  1
mm of H2SO4  = 5 mm.
Now H2SO4 that reacted with ammonia in actual is
=> 0.1 x 100 – 5
=> 10 –  5 
=> 5 mm.
The reaction is:
2NH3       +      H2SO4         --------->             (NH4)2SO4
From 5 mm of H2SO4 NH3 mmoles are => 5 x 2 =  > 10 mm.
mm of NH3 = 10 mm.
Applying POAC
For each option and cross checking,
Thaking Urea first,
for Nitrogen atom.
2 x moles = 10 x 1
moles =  5 mm.
For 0.3 gm,
MOlar masss = 0.3/0.005 => 60 g
Thus urea is the organic compound.